# Signature matrices of membranes

### v1

@@ 4.2 The case $mn\leq d$
The variety $\mathcal{P}_{d,m,n}$ is the Zariski closure of the image of the map $\phi:\mathbb{C}^{d\times mn}\to\mathbb{C}^{d\times d},\ S\mapsto SA_{m,n}S^{\top}$ where $A_{m,n}=\sigma^{(2)}(\text{Axis}^{m,n})$. Equivalently, it is the Zariski closure of the $\operatorname{GL}_{d}$-congruence orbit of the $d\times d$-matrix $A^{\prime}_{m,n}$ which is obtained from $A_{m,n}$ by extending with zeroes. This congruence orbit is a differentiable manifold and a semi-algebraic set, in particular the dimension of the manifold agrees with the dimension of its Zariski-closure. The dimension of the manifold agrees with the dimension of the tangent space at $\phi(I_{d})$ (as it is a group orbit). Its codimension is the dimension of the kernel of the Jacobian $J(\phi)$ at $I_{d}$, in other words, the dimension of the solution space to the matrix equation
$SA^{\prime}_{m,n}+A^{\prime}_{m,n}S^{\top}=0,$ (24)
cf. [45, Lemma 2]. This dimension remains unchanged when $A^{\prime}_{m,n}$ is replaced by $A_{m,n}$. Thus in the following, we may assume $d=mn$.
[…]
As $A_{m,n}$ is indeed non-singular (by Example 3.17 we have $\det(A_{m,n})=\frac{1}{4^{mn}}$), our strategy is thus to determine the Jordan canonical form of $A_{m,n}^{-\top}A_{m,n}$. This is simplified significantly by the following observation: if we denote $A_{u}:=\sigma^{(2)}(\mathsf{Axis}^{u})$ then by Proposition 3.3 we have $A_{m,n}=A_{m}\otimes A_{n}$ where $\otimes$ denotes the Kronecker product. In particular
$A_{m,n}^{-\top}A_{m,n}=(A_{m}\otimes A_{n})^{-\top}(A_{m}\otimes A_{n})=A_{m}^{-\top}A_{m}\otimes A_{n}^{-\top}A_{n}.$
Moreover, as observed in the proof of [2, Theorem 3.4], the matrix $A_{u}$ is congruent to $N_{u}:=\begin{pmatrix}1&1\\ -1&0\end{pmatrix}\oplus\tau^{\frac{u-2}{2}}$ if $u$ is even, and congruent to $N_{u}:=\begin{pmatrix}1\end{pmatrix}\oplus\tau^{\frac{u-1}{2}}$ if $u$ is odd, where $\tau=\begin{pmatrix}1&-1\\ -1&0\end{pmatrix}$. Thus, we may replace $A_{m,n}$ by $N_{m}\otimes N_{n}$, $A_{m}$ by $N_{m}$ and $A_{n}$ by $N_{n}$. These observations reduce the problem to the following two questions:
@@
    What are the Jordan canonical forms of $N_{u}^{-\top}N_{u}$?
@@
For the first question, note that $\tau^{-\top}\tau=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}=J_{1}(-1)$ and
The answer to the second question is [7, Theorem 4.6] from which we obtain: the Jordan canonical form of $A_{m,n}^{-T}A_{m,n}$ is
Thus the canonical form for congruence of $A_{m,n}$ is
Finally, we conclude by calculating the codimensions using the recipe given in [45, Theorem 2]. ∎
$\left[\begin{array}[]{cccc}4&7&9&10\\ 7&14&\mathbf{16}&16\\ 9&\mathbf{16}&\mathbf{16}&16\\ 10&16&16&16\end{array}\right]$
@@
$\left[\small\begin{array}[]{ccccc}5&9&12&14&15\\ 9&18&\mathbf{24}&\mathbf{25}&25\\ 12&\mathbf{24}&\mathbf{25}&\mathbf{25}&25\\ 14&\mathbf{25}&\mathbf{25}&\mathbf{25}&25\\ 15&25&25&25&25\end{array}\right]$
@@
$\left[\small\begin{array}[]{cccccc}6&11&15&18&20&21\\ 11&22&31&\mathbf{34}&\mathbf{36}&36\\ 15&31&\mathbf{34}&\mathbf{36}&\mathbf{36}&36\\ 18&\mathbf{34}&\mathbf{36}&\mathbf{36}&\mathbf{36}&36\\ 20&\mathbf{36}&\mathbf{36}&\mathbf{36}&\mathbf{36}&36\\ 21&36&36&36&36&36\end{array}\right]$
@@
$\left[\small\begin{array}[]{ccccccc}7&13&18&22&25&27&28\\ 13&26&37&\mathbf{43}&\mathbf{48}&\mathbf{49}&49\\ 18&37&\mathbf{43}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 22&\mathbf{43}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 25&\mathbf{48}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 27&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&\mathbf{49}&49\\ 28&49&49&49&49&49&49\end{array}\right]$
@@
$\left[\small\begin{array}[]{cccccccc}8&15&21&26&30&33&35&36\\ 15&30&43&52&\mathbf{60}&\mathbf{62}&\mathbf{64}&64\\ 21&43&\mathbf{52}&\mathbf{63}&\mathbf{63}&\mathbf{64}&\mathbf{64}&64\\ 26&52&\mathbf{63}&\mathbf{63}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 30&\mathbf{60}&\mathbf{63}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 33&\mathbf{62}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 35&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&\mathbf{64}&64\\ 36&64&64&64&64&64&64&64\\ \end{array}\right]$
@@
Figure 5: The dimensions $\dim(\mathcal{P}_{d,m,n})$ stored in $d\times d$ matrices for varying $d$, where bold numbers are not covered by Theorem 4.7 or Theorem 4.9. We compute these remaining dimensions according to Remark 4.13. Remarkable is the fact that in all examples the full dimension $d^{2}$ is already attained on the antidiagonal $m+n=d+1$ for $m,n\geq 2$, in particular $\mathcal{P}_{d,d-1,2}$ has full dimension, which is not clear from Remark 4.8.
@@
In light of Lemma 4.5, one checks that for $n=1$ the formulas simplify to $md-\binom{m}{2}$, which is indeed the dimension of the affine cone of the path signature matrix varieties considered in [2], cf. Theorem 3.4 in loc. cit..
When rescaling the rows of $\sigma(\mathsf{Mom}^{m})$ we obtain a Cauchy matrix, hence with Corollary 3.6, $\sigma(\mathsf{Mom}^{m,n})$ has full rank for any $m$ and $n$. However, the rank of the symmetric and skew-symmetric part behaves differently as in the path case:
Corollary 4.11.
For $A_{m,n}:=\sigma^{(2)}(\mathsf{Mom}^{m,n})$ let $A_{m,n}^{\mathsf{sym}}$ and $A_{m,n}^{\mathsf{sk}}$ denote its symmetric and skew-symmetric part, respectively. Then
$\operatorname{rk}(A_{m,n}^{\mathsf{sym}})=\begin{cases}m\cdot n&m,n\text{ even}\\ 1+m(n-1)&m\text{ even},n\text{ odd}\\ 1+(m-1)(n-1)&m,n\text{ odd}\end{cases}$
(again we omit the case $m$ odd, $n$ even due to symmetry) and
$\operatorname{rk}(A_{m,n}^{\mathsf{sk}})=\begin{cases}m+n-1&m+n\text{ odd}\\ m+n-2&m+n\text{ even.}\end{cases}$
Proof.
In the proof of Theorem 4.9, add up the ranks of the symmetric/skew-symmetric parts of the blocks in the canonical form for congruence of $A_{m,n}$. ∎
For $m=1$ or $n=1$ we obtain a rank one symmetric part according to the path framework, [2, eq. (4)].
As the rank of symmetric and skew-symmetric part is preserved under congruence, Corollary 4.11 induces constraints on the congruence orbits. However not all relations on the orbit arise in this way:
@@
Consider the case $d=4,m,n=2$. According to Corollary 4.11 the symmetric part of $A:=\sigma^{(2)}(\mathsf{Axis}^{m,n})$ has full rank and the skew-symmetric part has rank $2$. In particular, the Pfaffian $x_{12}x_{34}-x_{13}x_{24}+x_{14}x_{23}$ vanishes on the skew-symmetric part of matrices in the congruence orbit. However, according to Theorem 4.9 the dimension of $\mathcal{P}_{4,2,2}$ is 14, so this can not be the only relation. Indeed, another relation is given by $\det(A)\det(X^{\mathsf{sym}})=\det(X)\det(A^{\mathsf{sym}})$: the ratio $\det(X)/\det(X^{\mathsf{sym}})$ is well-defined and clearly constant on the congruence orbit. Using Macaulay 2 one checks that the variety $\mathcal{P}_{4,2,2}$ is a complete intersection, cut out by those two equations.
Remark 4.13.
Some choices of $m,n,d$ are neither covered by Theorem 4.7 nor by Theorem 4.9. Using Macaulay 2 with NumericalImplicitization, we computed the missing dimensions of $\mathcal{P}_{d,m,n}$ for $m,n\leq d$. We summarize our findings in Figure 5.
Example 4.14.
According to Figure 5, $\mathcal{P}_{5,3,2}$ is a hypersurface in $\mathbb{R}^{5\times 5}$. By numerical computations in Macaulay 2 using the packages NumericalImplicitization and MultigradedImplicitization one observes that this hypersurface is cut out by a single quintic $f$ with $999$ terms. One checks that $f\in\mathsf{minors}(3,X)\cap\mathsf{minors}(4,X^{\mathsf{sk}})\cap\mathsf{minors}(1,X^{\mathsf{sym}})$, where $\mathsf{minors}(i,X)$ denotes the ideal of $i$-minors of $X$.
5 Conclusions and future work
Our notion of signature was introduced by [26, 14] under the name of $\mathsf{id}$-signature, see Remark 2.3 for a detailed discussion. Both references observe that this $\mathsf{id}$-signature is not universal. We prove that in general the signature matrix of a membrane does not satisfy any algebraic relations, so in particular, no shuffle relations. This proof relies on matrix theory and is not generalizable for tensors in an obvious way. We conjecture that our result is true for all signature tensors.

### v2

@@ 4 Signature matrix varieties
This follows from Lemma 4.6 by using the codimension count in [TERAN201144, Theorem 2]. ∎
Remark 4.8.
In light of Lemma 4.5, one checks that for $n=1$ the polynomials in Corollary 4.7 simplify to $md-\binom{m}{2}$, which is indeed the dimension of the affine cone of the path signature matrix varieties considered in [bib:AFS2019, Theorem 3.4].
Corollary 4.9.
Let $\mathcal{S}_{a,b}$ be the variety of $d\times d$ matrices whose symmetric part has rank $\leq a$ and whose skew-symmetric part has rank $\leq b$. Then
@@
    $\mathcal{S}_{(m-2)(n-2),m+n-2}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{mn,m+n-2}$ if $m,n$ are even,
@@
    $\mathcal{S}_{2,n-1}\subseteq\mathcal{M}_{d,2,n}\subseteq\mathcal{S}_{2(n-1)+1,n+1}$ if $n$ is odd,
  3. 3.
    $\mathcal{S}_{(m-2)(n-1)-1,m+n-1}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{m(n-1)+1,m+n-1}$ if $m\geq 4$ is even, $n$ is odd,
  4. 4.
    $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}=\mathcal{M}_{d,m,n}$ if $m,n$ are odd.
Proof.
For the upper bounds, we just need to compute the ranks of the symmetric and skew-symmetric parts of the normal forms in Lemma 4.6.
  • For $m,n$ even the rank of $C_{m,n}^{\mathsf{sym}}$ is $(m-2)(n-2)+1+2m+2n-8+3=mn$. The rank of $C_{m,n}^{\mathsf{sk}}$ is $22+m+n-4=m+n-$ 2.
  • For $m$ even, $n$ odd the rank of $C_{m,n}^{\mathsf{sym}}$ is $11+2(n-1)+(m-2)(n-1)=m(n-1)+$ 1. The rank of $C_{m,n}^{\mathsf{sk}}$ is $12+n-1+m-2=m+n-$ 1.
  • For $m,n$ odd the rank of $C_{m,n}^{\mathsf{sym}}$ is $(m-1)(n-1)+1$. The rank of $C_{m,n}^{\mathsf{sk}}$ is $m+n-2$.
For the lower bound, note that for a general matrix $X\in\mathcal{S}_{a,b}$ there are $S,T\in\operatorname{GL}_{d}(\mathbb{C})$ with $S^{\top}X^{\mathsf{sym}}S=\Gamma_{1}^{\oplus a^{\prime}}$ and $T^{\top}X^{\mathsf{sk}}T=H_{2}(-1)^{\oplus b^{\prime}}$ where $a^{\prime}\leq a$ and $b^{\prime}\leq\frac{b}{2}$. Thus, to give a lower bound, one can count blocks in the normal form of $C_{m,n}$ that are congruent to $H_{2}(-1)$ and can thus contribute to the skew-symmetric rank, such that all other blocks are congruent to $\Gamma_{1}$ and can contribute to the symmetric rank.
For $m,n$ even, we have that $\Gamma_{3}\oplus\Gamma_{1}$ is congruent to $H_{2}(-1)$, as is $H_{4}(1)$. Thus, every general skew-symmetric matrix of rank $42+m+n-$ 4 is congruent to $\Gamma_{3}\oplus H_{4}(1)^{\oplus\frac{m+n-4}{2}}\oplus\Gamma_{1}$. As every rank $(m-2)(n-2)$ symmetric matrix is congruent to $\Gamma_{1}^{\oplus(m-2)(n-2)}$ we conclude.
For $m=2$, $n$ odd, we have that $\Gamma_{2}\oplus H_{4}(1)$ is congruent to $H_{2}(-1)\oplus\Gamma_{1}^{\oplus 2}$. Every general skew-symmetric matrix of rank $12+n-3=n-$ 1 is congruent to $H_{2}(-1)\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}$ and every symmetric matrix of rank $2$ is congruent to $\Gamma_{1}^{\oplus 2}$ and we conclude.
For $m$ even, $n$ odd, and $m\geq 4$ we know that $\Gamma_{2}\oplus\Gamma_{1}$ is congruent to $H_{2}(-1)$; thus every general skew-symmetric matrix of rank $12+n-1+m-2=m+n-$ 1 is congruent to $\Gamma_{2}\oplus H_{4}(1)^{\oplus\frac{n-1}{2}}\oplus H_{2}(-1)^{\oplus\frac{m-2}{2}}\oplus\Gamma_{1}$; and we conclude again.
By the same reasoning, we immediately see that for $m,n$ odd we just obtain the variety $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}$ as a lower bound as well and deduce equality with $\mathcal{M}_{d,m,n}$. ∎
Note that for $m=1$ or $n=1$ we obtain a rank one symmetric part in accordance with the path framework, [bib:AFS2019, eq. (4)]. Moreover, the corollary shows for all cases except $m=2$, $n$ odd, that the projection of matrices in $\mathcal{M}_{d,m,n}$ to the skew-symmetric part gives precisely the variety of skew-symmetric matrices of rank $\leq m+n-1$ (if $m+n$ is odd) or $\leq m+n-2$ (if $m+n$ is even).
In the case where $m,n$ are odd, we can describe the variety precisely:
Corollary 4.10.
The degree of $\mathcal{M}_{d,m,n}$ is
$\frac{1}{2^{d-(m+n)+2}}\prod_{\alpha=0}^{d-(m-1)(n-1)}\frac{\binom{d+\alpha}{d-(m-1)(n-1)+1-\alpha}}{\binom{2\alpha+1}{\alpha}}\prod_{\alpha=0}^{d-(m+n)}\frac{\binom{d+\alpha}{d-(m+n)+1-\alpha}}{\binom{2\alpha+1}{\alpha}}.$
Proof.
This follows from Corollary 4.9 since $\mathcal{M}_{d,m,n}=\mathcal{S}_{(m-1)(n-1)+1,m+n-2}$. Indeed, for $m,n$ odd, $m+n\leq d$, $\mathcal{S}_{(m-1)(n-1)+1,m+n-2}$ is the transverse intersection of the vanishing locus of the $((m-1)(n-1)+2)$-minors of the symmetric part with the vanishing locus of the $(m+n)$-Pfaffians of the skew-symmetric part. The degree of these two vanishing loci is calculated in [HARRIS198471, Proposition 12]. ∎
Example 4.11.
According to Corollary 4.7 and Corollary 4.10 the variety $\mathcal{M}_{6,3,3}$ has dimension $34$ and degree $18$. It is a complete intersection of the vanishing locus of the Pfaffian of the skew-symmetric part and the vanishing locus of the determinant of the symmetric part.
Example 4.12.
Consider the case $d=4$ and $m=n=2$. By Corollary 4.9, we have $\mathcal{S}_{0,2}\subseteq\mathcal{M}_{d,m,n}\subseteq\mathcal{S}_{4,2}$. In particular, the Pfaffian $x_{12}x_{34}-x_{13}x_{24}+x_{14}x_{23}$ vanishes on the skew-symmetric part of matrices in the congruence orbit. However, according to Corollary 4.7 the dimension of $\mathcal{M}_{4,2,2}$ is 14, so there must be additional relations on the symmetric part. Indeed, another relation is given by $\det(A)\det(X^{\mathsf{sym}})=\det(X)\det(A^{\mathsf{sym}})$: the ratio $\det(X)/\det(X^{\mathsf{sym}})$ is well-defined and clearly constant on the congruence orbit. Using Macaulay 2 one checks that the variety $\mathcal{M}_{4,2,2}$ is a complete intersection, cut out by those two equations.
It is immediate from Corollary 4.9 that $\mathcal{M}_{d,m,n}$ attains the full ambient dimension $d^{2}$ when $m,n\to\infty$. Our main theorem gives a precise bound:
Theorem 4.13.
For $m,n\geq 2$, we have $\mathcal{M}_{d,m,n}=\mathbb{C}^{d\times d}$ whenever $m+n\geq d+1$.
Proof.
Clearly, it suffices to prove the statement for $d=m+n-1$. We make a case distinction for the different parities of $m$ and $n$. If $m$ and $n$ are odd, then we have $(m-1)(n-1)+1\geq m+n-1$. Moreover, $d=m+n-1$ is odd, so that every $d\times d$ matrix has skew-symmetric rank $\leq m+n-2$. Thus, we immediately conclude by Corollary 4.9.
We prove the statement for the remaining cases by induction on $m+n$. We settle the base cases by explicit linear algebra calculations.
The base case. Consider the map $\phi:\mathbb{C}^{mn\times d}\to\mathbb{C}^{d\times d}$ given by $A\mapsto A^{\top}C_{m,n}A$. Let us write $\phi_{B}^{\prime}$ for its derivative
$\mathbb{C}^{mn\times d}\to\mathbb{C}^{d\times d},A\mapsto B^{\top}C_{m,n}A+A^{\top}C_{m,n}B$
at a matrix $B\in\mathbb{C}^{mn\times d}$. It suffices to give a matrix $B$ where $\phi_{B}^{\prime}$ has full rank.
The base case for $m,n$ even is $m=n=2$, $d=3$. Then the canonical form of $A_{2,2}$ is $\Gamma_{3}\oplus J_{1}(1)$. We compute that for $B=\begin{pmatrix}\mathrm{I}_{d}&\mathbbm{1}\end{pmatrix}$, $\phi_{B}^{\prime}$ has full rank.
The base case for $m$ even, $n$ odd is $m=2$, $n=3$, $d=4$. We compute that for $B=\begin{pmatrix}\mathrm{I}_{d}&\mathbbm{1}&\mathbbm{1}\end{pmatrix}$, $\phi_{B}^{\prime}$ has full rank.
The induction step. Assume the statement is true for $m+n=d+1$ for some $d$. We claim that it then also holds for $m+n=d+3=(d+2)+1$, i.e.  $\mathcal{M}_{d+2,m,n}=\mathbb{C}^{d+2,d+2}$. By assumption, we have $\mathcal{M}_{d,m,n-2}=\mathbb{C}^{d\times d}$ or $\mathcal{M}_{d,m-2,n}=\mathbb{C}^{d\times d}$. A general $\mathbb{(}d+2)\times(d+2)$-matrix has normal form $\oplus_{i\leq k}H_{2}(\mu_{i})\oplus\Gamma_{1}^{\oplus l}$ where $22k+l=d+$ 2 with $l\in\{0,1\}$ and $\mu_{i}\in\mathbb{C}$. In particular, wlog. we can replace a general $(d+2)\times(d+2)$ matrix by a matrix $D\oplus E$ for $D$ a general $d\times d$-matrix and $E$ a general $2\times 2$-matrix.
If $m$ and $n$ are even and wlog. $m\geq 4$, then we have $C_{m,n}=C_{m-2,n}\oplus H_{4}(1)\oplus\Gamma_{1}^{\oplus 2(n-2)+1}$. A calculation in Macaulay 2 shows that generically the fibers of the map $\mathbb{C}^{4\times 2}\to\mathbb{C}^{2\times 2},A\mapsto A^{\top}H_{4}(1)A$ are $5$-dimensional.
If $m$ is even and $n$ is odd, then if $m\geq 4$ we have $C_{m,n}=C_{m-2,n}\oplus H_{2}(-1)\oplus\Gamma_{1}^{\oplus 2(n-1)}$. Again, a calculation in Macaulay 2 shows that generically the fibers of the map $\mathbb{C}^{4\times 2}\to\mathbb{C}^{2\times 2},A\mapsto A^{\top}(H_{2}(-1)\oplus\Gamma_{1}^{\oplus 2})A$ are $5$-dimensional. If $n\geq 5$, then $C_{m,n}=A_{m,n-2}\oplus H_{4}(1)\oplus\Gamma_{1}^{\oplus 2(m-2)}$ and we conclude as in the case $m,n$ even.
Finally, note that for matrices $A,B,X,Y$ we have $(A\oplus B)^{\top}(X\oplus Y)(A\oplus B)=A^{\top}XA\oplus B^{\top}YB$. Thus by the induction hypothesis we conclude in all cases that $\mathcal{M}_{d+2,m,n}=\mathbb{C}^{d+2,d+2}$. ∎
For $m=1$ or $n=1$ we obtain a rank one symmetric part according to the path framework, [bib:AFS2019, eq. (4)].
$\left[\begin{array}[]{cccc}4&7&9&10\\ 7&14&16&\\ 9&16&&\\ 10&&&\end{array}\right]$
@@
$\left[\small\begin{array}[]{ccccc}5&9&12&14&15\\ 9&18&\mathbf{24}&25&\\ 12&\mathbf{24}&25&&\\ 14&25&&&\\ 15&&&&\end{array}\right]$
@@
$\left[\small\begin{array}[]{cccccc}6&11&15&18&20&21\\ 11&22&31&\mathbf{34}&36&\\ 15&31&\mathbf{34}&36&&\\ 18&\mathbf{34}&36&&&\\ 20&36&&&&\\ 21&&&&&\end{array}\right]$
@@
$\left[\small\begin{array}[]{ccccccc}7&13&18&22&25&27&28\\ 13&26&37&\mathbf{43}&\mathbf{48}&49&\\ 18&37&\mathbf{43}&\mathbf{49}&&&\\ 22&\mathbf{43}&\mathbf{49}&&&&\\ 25&\mathbf{48}&&&&&\\ 27&49&&&&&\\ 28&&&&&&\end{array}\right]$
@@
$\left[\small\begin{array}[]{cccccccc}8&15&21&26&30&33&35&36\\ 15&30&43&52&\mathbf{60}&\mathbf{62}&64&\\ 21&43&\mathbf{52}&\mathbf{63}&\mathbf{63}&64&&\\ 26&52&\mathbf{63}&\mathbf{63}&64&&&\\ 30&\mathbf{60}&\mathbf{63}&64&&&&\\ 33&\mathbf{62}&64&&&&&\\ 35&64&&&&&&\\ 36&&&&&&&\\ \end{array}\right]$
@@
Figure 6: The dimensions $\dim(\mathcal{M}_{d,m,n})$ stored in $d\times d$ matrices for varying $d$, where bold numbers are not covered by Theorem 4.13 or Corollary 4.7. We omit the values once the full dimension has been reached. We compute these remaining dimensions according to Remark 4.14.
@@
Some choices of $m,n,d$ are neither covered by Theorem 4.13 nor by Corollary 4.7. Using Macaulay 2 with NumericalImplicitization [chen2019numerical], we computed the missing dimensions of $\mathcal{M}_{d,m,n}$ for $m,n\leq d$. We summarize our findings in Figure 6.
@@
Due to Figure 6, $\mathcal{M}_{5,3,2}$ is a hypersurface in $\mathbb{R}^{5\times 5}$. Using the packages NumericalImplicitization and MultigradedImplicitization [cummings2025computing], one observes that this hypersurface is cut out by a single quintic $f$ with $999$ terms. One checks that $f\in\mathsf{minors}(3,X)\cap\mathsf{minors}(4,X^{\mathsf{sk}})\cap\mathsf{minors}(1,X^{\mathsf{sym}})$, where $\mathsf{minors}(i,X)$ denotes the ideal of $i$-minors of $X$ from the congruence orbit.
5 Outlook and future work
Our notion of signature was introduced by [GLNO2022, diehl2024signature] under the name of $\mathsf{id}$-signature, see Remark 2.3 for a detailed discussion. Both references observe that this $\mathsf{id}$-signature is not universal. We prove that in general the signature matrix of a membrane does not satisfy any algebraic relations, so in particular, no shuffle relations. This proof relies on matrix theory and is not generalizable for tensors in an obvious way. We conjecture that our result is true for all signature tensors.